Assume that a 5-lb ideal gas with 𝑅 = 38.7 𝑓𝑑.𝑙𝑏 𝑙𝑏.°𝑅 and π‘˜ = 1.668 have 300 Btu of heat added during reversible constant pressure change of state. The initial temperature is 80℉. Determine (a) the final temperature; (b) the change of enthalpy, internal energy, and entropy; and (c) the work.


1
Expert's answer
2021-10-28T08:51:00-0400

(a)

"m=5 \;lb \\

R = 38.7 \;ft \cdot lb/lbm \cdot R \\

k = 1.668 \\

Q = 300 \;BTU = 233451 lbf \cdot ft \\

T_1 = 80 \\

Q = mc_p(T_2-T_1) \\

233451 = 5 \times \frac{kR}{k-1}(T_2-T_1) \\

233451 = 5 \times \frac{1.668 \times 38.7}{0.668} \times (T_2-80) \\

T_2 = 563.165 \; °F"

(b)

"ΔU = mc_v(T_2-T_1) \\

= \frac{mR}{k-1}(T_2-T_1) \\

= 5 \times \frac{38.7}{0.668} \times (563.165-80) \\

= 139958.6331 \; lbf \cdot ft"

(c)

"ΔH = Q = 233451 \;lbf \cdot ft \\

ΔS = mc_pln(\frac{T_2}{T_1}) \\

= 5 \times \frac{1.668}{0.668} \times 38.7 \times ln(\frac{563.165}{80}) \\

= 942.9295"

(d)

"W = Q -ΔU \\

= 233451 -139958.633 \\

= 93492.3669 \; lbf \cdot ft"